Question 23 . > 1 if x = 3 and f(x) = x2 – 3 and g(x) = then g(f(x))Find the following given: f(x) = x2 + 3 and g(x) = x – 2 f(g(x))= Submit Question
Answer #1

g(f(0) and  \large f(g(x)) are known as composition function. To solve for g(f(0) , first write the given value of \large f(x)   , then express  \large f(x) as x is expressed in the function \large g(x) .

Similarly we can do the same for solving \large f(g(x)) .

From algebra, \large (a-b)^2=a^2-2(a)b)+b^2

  \large f(x)=x^2-3 and  \large g(x)=\frac{1}{x}

then \large g(f(x))=   \large g(x^2-3)=  \large \frac{1}{x^2-3}  

So,   \large g(f(x))=\frac{1}{x^2-3}

When,  x=3, then ,

  \large g(f(3)) = \frac{1}{3^2-3}=\frac{1}{9-3}=\frac{1}{6}

Now, \large f(x)=x^2+3 and   \large g(x)=x-2

  \large f(g(x))=f(x-2) = (x-2)^2+3

So,  \large f(g(x))=(x^2-2(x)(2)+2^2)+3

  \large =(x^2-4x+4)+3

  \large =x^2-4x+7

  \large f(g(x))=x^2-4x+7

When x=3,

\large f(g(3))=3^2-4(3)+7

\large =9-12+7

\large =4